OpenStudy (anonymous):

Write the standard form of the equation of the circle with the given characteristics center (3,-2) solution point (-1,1)

OpenStudy (zale101):

standard equation of a circle. (x - h)^2 + (y - k)^2 = r^2 r = [(x2-x1)^2 + (y2-y1)^2]^1/2 r = [(-1 + 3)^2 + (1+2)^2]^1/2